OpenStudy (anonymous):

Use the Gauss Jordan elimination method to find all solutions of the system of equations: WILL GIVE A MEDAL!!

OpenStudy (anonymous):

\[2x _{1}+3x _{2}=12\]

OpenStudy (anonymous):

\[-4x _{1} + 4x _{2} = -4\]

OpenStudy (anonymous):

\[8x _{1}-23x _{2} = -22\]

OpenStudy (anonymous):

Find all solutions of the system of equations.

OpenStudy (anonymous):

@genius12

hero (hero):

Original System\[2x_1 + 3x_2 = 12 \\-4x_1 + 4x_2 = -4 \\8x_1 - 23x_2 = -22\] Divide R2 by -4: \[2x_1 + 3x_2 = 12 \\x_1 - x_2 = 1 \\8x_1 - 23x_2 = -22\] Swap R2 and R1: \[x_1 - x_2 = 1 \\2x_1 + 3x_2 = 12 \\8x_1 - 23x_2 = -22\] Add -2R1 to R2: \[x_1 - x_2 = 1 \\5x_2 = 10 \\8x_1 - 23x_2 = -22\] Divide R2 to 5: \[x_1 - x_2 = 1 \\x_2 = 2 \\8x_1 - 23x_2 = -22\] Add R2 to R1: \[x_1 = 3 \\x_2 = 2 \\8x_1 - 23x_2 = -22\] Add 23*R2 to R3: \[x_1 = 3 \\x_2 = 2 \\8x_1 = 24\] Add -8*R1 to R3: \[x_1 = 3 \\x_2 = 2\]

OpenStudy (anonymous):

We still start off by making an Augmented matrix then apply elemental row operations to solve for A. We will call our matrix. NOTE: In this first matrix, the reason for why you don't see the third row is because I multiplied row 3 by 0 so that it would be eliminated. The reason I eliminated it by multiplying it by 0 is because we only have 2 variables to solve for; x1 and x2. The first two rows representing the 2 equations are enough to solve for x1 and x2 so I thought it would be best to eliminate row 3 first. \[A=\left[\begin{matrix}2 & 3 & 12 \\ -4 & 4 & -4\end{matrix}\right]\]Now we apply two elemental row operations to this matrix: row2 + (2 * row1) and row1 - (row2 * 3/4). This gives us the following matrix:\[A=\left[\begin{matrix}5 & 0 & 15 \\ 0 & 10 & 20\end{matrix}\right]\]In Gauss-Jordan Elimination, we need to have a diagonal of 1s and the rest be 0s. Thus I apply 2 more operations: row1 / 5 and row / 10. This gives us our final 'Identity Matrix':\[A=\left[\begin{matrix}1 & 0 & 3 \\ 0 & 1 & 2\end{matrix}\right]\]From this matrix, the values in the third column are basically our solutions to x1 and x2. \[\therefore x_{1}=3,x _{2}=2\]@Anita505 Sorry for the wait. I know Hero has already answered but I solved the matrix in a much faster way with only 2 steps, so I thought this might help you out.

OpenStudy (anonymous):

@Anita505

OpenStudy (anonymous):

I know lol. But I eliminated the third row from the matrix by multiplying it by 0 in order to get the same solution. And yes I know I can't ignore the third one, but I don't know how to add a third row with the equation box so I just eliminated it lol. And this is case, you should get the same solution with or without the third row which is why I ignored it.

OpenStudy (anonymous):

And btw, I didn't start this question till much later on =.=

hero (hero):

Multiplying the third row by zero is an illegal elementary operation.

hero (hero):

So you did not solve it properly. Sorry bro.

hero (hero):

If you were a student and I were the teacher, I would mark it wrong.

OpenStudy (anonymous):

I know that multiplying by 0 is illegal, but the point of saying that was just that you don't need the third to solve for this system. It's the same thing without it. That's what I meant by 'multiplying by 0'.

hero (hero):

I understand what you're trying to say, but your steps (no matter how you try to justify them) are incorrect according to the rules for elementary operations.

OpenStudy (anonymous):

The justification for why you can eliminate the third row in this case is that if one can solve for the variables of a system without involving a certain row in the elemental row operations, then that row can be eliminated from the system as it isn't required to solve for the variables.

hero (hero):

You said you multiplied the third row by 0 to eliminate it. That is not correct.

OpenStudy (anonymous):

I did. But I have provided a justification which should be enough.

hero (hero):

Well, the problem is, what you did was not Gauss-Jordan Elimination

OpenStudy (anonymous):

Forget it....=.=